Assume that the economy is$500 billion above its full-employment GDP and the marginal propensity to consume is 0.75 Give a fiscal policy recommendation to close the gap in GDP. How much is needed? Note: This will depend on what fiscal policy tool you recommend in a) Show your calculations to receive full points. (you can assume AD shortfall = GDP gap) How would this policy change affect the economy's budget position?
Answer #1

As the economy is above the equilibrium level by $500 billion the government in the market needs to cut the expenditure to bring the economy back at the equilibrium level.

And as the MPC is 0.75 the multiplier in the market will be 1 / 1 - MPC = 1 / 1- 0.75 = 4.

As the multiplier in the market is 4 to reduce the output in the market by 500 billion the government need to reduce the expenditure by 500 /4 = 125 billion.

this policy will reduce the budget deficit in the market.